The following cell has a potential of +1.2272 V at 25°C: Zn(s)|ZnCl2(aq 0.0005m)||Hg2Cl2(s)|Hg(l)|Pt(s). Determine E for this cell assuming the Debye Hückel equation is satisfied.



1
Expert's answer
2022-01-10T14:01:42-0500



Zn(s)"+" Hg2+(aq) "\rightleftharpoons\>" "Zn^{2+}_{(aq)}+Hg_{(l)}"


"E=E^0\>'+\frac{RT}{nF}" In "\frac{C^*_0}{C^*_R}"



Where "E^0\>'" (V) is formal potential


"C^*[Mol\>L^{-1}]" is the bulk concentration for the considered species


"\frac{RT}{nF}=\frac{8.314×298.15}{2×96485}=0.01285"


The activity for a very dilute solution is the same as the concentration


"\therefore\>E=1.2272+0.01285\> 
In\> \>\frac{0.0005}{1}"


"=1.1295V"